### v1

@@ 1 Introduction
Thus, previous works [14, 15, 8] have tried to define coherence and repetitiveness metrics and used them to measure the ability of story generators. However, their attempts are also limited only to short sentence generation problems and have never considered completeness. Completeness is the ability to conclude a story of any length properly. It is a significant metric not only for evaluating story generation but also for a wide range of open-domain generation and dialogue system tasks. Therefore, in our model evaluation, we consider not only coherence and repetitiveness but also prioritize completeness as a critical measure of performance.
In this paper, we tackle this challenging problem by introducing a novel long story generator, Coherent, Complete and Length Controlled Long story Generator(LongStory), which covers from a few hundred tokens to 10k tokens length, not limited to only a few sentences. The LongStory utilizes two novel methods: (1) long and short-term contexts weight calibrator (CWC) and (2) long story structural positions (LSP) to tackle coherence and completeness. Specifically, we implement the CWC using BERT-tiny to calibrate the weight to which long-term and short-term contexts are utilized since both contexts contribute differently in every paragraph when writing a story. For the LSP, we use discourse tokens, which means the order of paragraphs, to give information about the structural positions of paragraphs to the story generator (e.g., $<front>$, $<middle>$,$<ending>$, $<next\_is\_ending>$). Our model uses more abundant discourse tokens than the previous study [8] since a detailed understanding of story structure is essential to generate much longer posts.
@@ 4.1 Experiments Set-up
We train our models on Writing Prompts[4], Booksum[16], and Reedsy Prompts [^3] We crawl this dataset from https://blog.reedsy.com/short-stories/. datasets. For Booksum, we use the ‘Chapter’ column as a dataset. Since our task is to generate various lengths of stories with coherence and completeness, we must prepare datasets of different lengths. We use the same train/valid/test split from the original papers[4, 16], but since Reedsy Prompts is not used in any other works, we split it into train/valid/test sets of 60k/4k/4k. The NLTK library’s sentence tokenizer is utilized for document segmentation into sentences during truncation, with each sentence sequentially compiled into a paragraph, ensuring a maximum length of 200 tokens per paragraph. For each dataset, we extract keywords using the RAKE algorithm[13] to make keyword-story pairs. These keywords are used as input for the model to write each paragraph of a story.
@@
We mainly compare our model to Plotmachine, the only model that performs paragraph-level recursive story generation. We also compare the performance of our model against GPT-2 and BART with the recursive generation setting, no applied memory, and cheating contexts; thus, it does not utilize the CWC. Plotmachine receives only the discourse tokens from the original paper, while other baselines are given the same as ours. We also compare our model with the ablated version. The LongStory with no memory (LongStory¬${}^{\neg M}$), LongStory with no cheating (LongStory¬${}^{\neg C}$) and LongStory with no new discourse (LongStory¬${}^{\neg D}$) show the effectiveness of context vectors and newly added discourse tokens.
@@ 4.2 Experimental Results
For the Coherence scorer, we divide two consecutive paragraphs into true pairs and two randomly selected paragraphs into fake pairs. We then train GPT-2 to use these constructed datasets on a classification task to distinguish whether the two paragraphs are true or fake folds[14, 20], and convert the result of it into a score $\in[0,1]$ using sigmoid function, as a Coherence score. We do the same for Completeness scorer, except for defining the true fold with the last paragraphs and the fake fold with the non-last paragraphs. It’s important to highlight that the reported Completeness scores are calculated as the last paragraph’s Completeness score minus the average Completeness scores of non-last paragraphs. This approach is chosen because an ideal story should generate a last paragraph appropriately and treat the body distinctly. Regarding Coherence, as shown in table  2, our model is the best or second best in the 5,10,19,30 paragraph length sample. Plotmachine had a higher Coherence score than BART but lower than our model or GPT-2. In the ablated analysis, LongStory¬${}^{\neg C}$ almost always outperforms LongStory¬${}^{\neg M}$ and LongStory¬${}^{\neg D}$ on Coherence. This suggests that $Memory$ is central to capturing the natural flow between paragraphs. In the Completeness, our model is the best for all lengths. LongStory¬${}^{\neg M}$ and LongStory¬${}^{\neg C}$ have no absolute advantage in this column, so it is difficult to know exactly how $Memory$ and $Cheating$ affect Completeness. Still, since our model performs the best, we can say that a balanced mix of the two contexts improves performance. The results for LongStory¬${}^{\neg D}$ also suggest that the newly added discourse tokens help improve Completeness.
@@
We also show the average ROUGE scores between the generated documents and the golden labels to evaluate how well the model reproduces the golden label from the keywords. This serves as a metric to assess the relevance of the keywords to the predictions. As shown in table  2, our model is best for results with 50 paragraph lengths and second best in many cases. Our model outperforms Plotmachine, GPT-2, and BART for all lengths of results. In the ablated analysis, LongStory¬${}^{\neg D}$ scores the best, for the most part. This shows that the newly added order tokens have little or even a negative effect on relevance. However, in the repeatability test in the next section, LongStory¬${}^{\neg D}$ performed worse than our model, which suggests that it reproduced the overall repetitive n-gram. For the most part, LongStory¬${}^{\neg M}$ performed better than LongStory¬${}^{\neg C}$, which means that $Cheating$ was more effective than $Memory$ in relevance.
@@
To see how much repetition occurs within a single output, we calculate the n-gram BLEU score by taking one paragraph as a hypothesis and the rest as a reference and averaging the results. We call the averaged results the in-SELF BLEU score [^5] Note that in-self-BLEU score is not the same as self-BLEU score[8]. The self-BLEU score has taken one whole generated document as a hypothesis and the others as references, which cannot represent inner repetitiveness.. The higher in-SELF BLEU score means that there are greater repetitions within the text and less diversity. As shown in table  2, Our model performs best on the in-self-BLEU score for all but five paragraphs length. For five paragraph lengths, Plotmachine performs best, and ours is second highest. However, for longer results, our model is always better. In the ablated analysis, LongStory¬${}^{\neg M}$ outperforms LongStory¬${}^{\neg C}$ and BART. This suggests that $Cheating$ prevented the model from over-attending to past context or repeating n-grams irrelevant to the overall context. The better performance of our model over LongStory¬${}^{\neg D}$ suggests that the order tokens we added positively affected reducing repetition.

### v2

@@ 1 Introduction
Thus, previous works [12, 13, 6] have tried to define coherence and repetitiveness metrics and used them to measure the ability of story generators.
However, their attempts are also limited only to short sentence generation problems and have never considered completeness. Completeness is the ability to conclude a story of any length properly. It is a significant metric not only for evaluating story generation but also for a wide range of open-domain generation and dialogue system tasks. Therefore, in our model evaluation, we consider not only coherence and repetitiveness but also prioritize completeness as a critical measure of performance.
In this paper, we tackle this challenging problem by introducing a novel long story generator, Coherent, Complete and Length Controlled Long story Generator(LongStory), which covers from a few hundred tokens to 10k tokens length, not limited to only a few sentences. The LongStory utilizes two novel methods: (1) long and short-term contexts weight calibrator (CWC) and (2) long story structural positions (LSP) to tackle coherence and completeness.
Specifically, we implement the CWC using BERT-tiny to calibrate the weight to which long-term and short-term contexts are utilized since both contexts contribute differently in every paragraph when writing a story. For the LSP, we use discourse tokens, which means the order of paragraphs, to give information about the structural positions of paragraphs to the story generator (e.g., $<front>$, $<middle>$,$<ending>$, $<next\_is\_ending>$). Our model uses more abundant discourse tokens than the previous study [6] since a detailed understanding of story structure is essential to generate much longer posts.
@@ 4.1 Experiments Set-up
We train our models on Writing Prompts[3], Booksum[14], and Reedsy Prompts [^3] For more information, https://huggingface.co/datasets/Iyan/reedsyPrompts. datasets. For Booksum, we use the ‘Chapter’ column as a dataset. Since our task is to generate various lengths of stories with coherence and completeness, we must prepare datasets of different lengths. We use the same train/valid/test split from the original papers[3, 14], but since Reedsy Prompts is not used in any other works, we split it into train/valid/test sets of 60k/4k/4k. The NLTK library’s sentence tokenizer is utilized for document segmentation into sentences during truncation, with each sentence sequentially compiled into a paragraph, ensuring a maximum length of 200 tokens per paragraph. For each dataset, we extract keywords using the RAKE algorithm[11] to make keyword-story pairs. These keywords are used as input for the model to write each paragraph of a story.
@@
We mainly compare our model to Plotmachine, the only model that performs paragraph-level recursive story generation. We also compare the performance of our model against GPT-2[28] and BART with the recursive generation setting, no applied memory, and cheating contexts; thus, it does not utilize the CWC. Plotmachine receives only the discourse tokens from the original paper, while other baselines are given the same as ours. We also compare our model with the ablated version. The LongStory with no memory (LongStory¬M), LongStory with no cheating (LongStory¬C) and LongStory with no new discourse (LongStory¬D) show the effectiveness of context vectors and newly added discourse tokens.
@@ 4.2 Experimental Results
For the Coherence scorer, we divide two consecutive paragraphs into true pairs and two randomly selected paragraphs into fake pairs. We then train GPT-2 to use these constructed datasets on a classification task to distinguish whether the two paragraphs are true or fake folds[12, 18], and convert the result of it into a score $\in[0,1]$ using sigmoid function, as a Coherence score. We do the same for Completeness scorer, except for defining the true fold with the last paragraphs and the fake fold with the non-last paragraphs. It’s important to highlight that the reported Completeness scores are calculated as the last paragraph’s Completeness score minus the average Completeness scores of non-last paragraphs. This approach is chosen because an ideal story should generate a last paragraph appropriately and treat the body distinctly. Regarding Coherence, as shown in table  2, our model is the best or second best in the 5,10,19,30 paragraph length sample. Plotmachine had a higher Coherence score than BART but lower than our model or GPT-2. In the ablated analysis, LongStory¬C almost always outperforms LongStory¬M and LongStory¬D on Coherence. This suggests that $Memory$ is central to capturing the natural flow between paragraphs. In the Completeness, our model is the best for all lengths. LongStory¬M and LongStory¬C have no absolute advantage in this column, so it is difficult to know exactly how $Memory$ and $Cheating$ affect Completeness. Still, since our model performs the best, we can say that a balanced mix of the two contexts improves performance. The results for LongStory¬D also suggest that the newly added discourse tokens help improve Completeness.
@@
We also show the average ROUGE scores between the generated documents and the golden labels to evaluate how well the model reproduces the golden label from the keywords. This serves as a metric to assess the relevance of the keywords to the predictions. As shown in table  2, our model is best for results with 50 paragraph lengths and second best in many cases. Our model outperforms Plotmachine, GPT-2, and BART for all lengths of results. In the ablated analysis, LongStory¬D scores the best, for the most part. This shows that the newly added order tokens have little or even a negative effect on relevance. However, in the repeatability test in the next section, LongStory¬D performed worse than our model, which suggests that it reproduced the overall repetitive n-gram. For the most part, LongStory¬M performed better than LongStory¬C, which means that $Cheating$ was more effective than $Memory$ in relevance.
@@
To see how much repetition occurs within a single output, we calculate the n-gram BLEU score by taking one paragraph as a hypothesis and the rest as a reference and averaging the results. We call the averaged results the in-SELF BLEU score [^5] Note that in-self-BLEU score is not the same as self-BLEU score[6]. The self-BLEU score has taken one whole generated document as a hypothesis and the others as references, which cannot represent inner repetitiveness.. The higher in-SELF BLEU score means that there are greater repetitions within the text and less diversity. As shown in table  2, Our model performs best on the in-self-BLEU score for all but five paragraphs length. For five paragraph lengths, Plotmachine performs best, and ours is second highest. However, for longer results, our model is always better. In the ablated analysis, LongStory¬M outperforms LongStory¬C and BART. This suggests that $Cheating$ prevented the model from over-attending to past context or repeating n-grams irrelevant to the overall context. The better performance of our model over LongStory¬D suggests that the order tokens we added positively affected reducing repetition.
@@ 5 Conclusion
6 Acknowledgements
We thank anonymous reviewers for their constructive and insightful comments. K. Jung is with ASRI, Seoul National University, Korea. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.2022-0-00184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics & NO.2021-0-01343, Artificial Intelligence Graduate School Program (Seoul National University)]